OpenStudy (anonymous):

Can some help me understand how i would go about solving a question like 3sqrt56^2

OpenStudy (anonymous):

Do you need to simplify? 3 sqrt 56^2 is just a number, but you can simplify...

OpenStudy (anonymous):

Thats the thing i'm not sure how to simplify this

OpenStudy (anonymous):

Just to be clear, how much is under the square root? Is it like:\[3\sqrt{56^{2}}\] or \[(3\sqrt{56})^{2}\] or something else? You can draw it if it's easier than using the equation tool.

OpenStudy (anonymous):

>.< Sorry i typed it wrong it's 15^14 and its in the form of the first one.

OpenStudy (anonymous):

-_- im still stupid 56^14

OpenStudy (anonymous):

So,\[3 \times \sqrt{56^{14}}\]

OpenStudy (anonymous):

Yes, it is

OpenStudy (anonymous):

Ok, good... just wanted to make sure we were talking about the same thing :) To simplify, notice that under the square root, 56^14 is like 56 * 56 * 56. * etc... right, fourteen repeats of multiplying the 56's? so it could be thought as seven pairs of 56's multiplied... (56*56) * (56*56) * (56*56) * (56*56) * (56*56) * (56*56) * (56*56) Does that make sense so far?

OpenStudy (anonymous):

Yes Sir!

OpenStudy (anonymous):

Great :) For each one of those (56*56) terms, when you take the square root, the result is 56. you have (56*56)^7 under the square root, so for each of those seven terms, when you simplify, 56 comes out from the square root.

OpenStudy (anonymous):

I'm with you still.

OpenStudy (anonymous):

Ok good